OpenStudy (anonymous):

Find an equation of the line that satisfies the given conditions. Through (1, −4); parallel to the line x + 2y = 6

OpenStudy (anonymous):

Since it is parallel to x+2y=6, this means that the slope of your new line remains the same. Thus, it will be of the form x+2y=k for some number k. Since you know that the line goes through (1,-4), plug these values in to get 1+2(-4)=-7, so k=-7. thus, the equation of the line is x+2y=-7